OpenStudy (anonymous):

pls help ..integrate (tan^-1y)/(1+y^2)

OpenStudy (dumbcow):

use u substitution u = tan^-1(y) du = 1/(1+y^2) dy --> dy = (1+y^2)du = integral u du = 1/2 u^2 + C =1/2(tan^-1(y))^2 + C

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

no prob